OpenStudy (anonymous):

A little boy is pulling a wagon full of 10 bricks. The mass of the wagon is too small to be considered. If the boy later is pulling the wagon with the same force and the wagon has 45 bricks in it, then how does the acceleration of the 45 brick wagon compare to the acceleration of the 10 brick wagon?

OpenStudy (anonymous):

let mass of 1 brick is m and force applied is f initially f=10*m*a1 a1=f/(10*m)...............(1) similarly for 45 bricks a2=f/(45*m)..............(2) now (1)/(2) a1/a2=4.5 a1=4.5*a2